in 4 years Daniel's age will be the same as Justin's age now. In 2 years,Justin will be twice as old as Daniel. Find their ages now.
Answer by ankor@dixie-net.com(22740)

In 4 years Daniel's age will be the same as Justin's age now. In 2 years, Justin will be twice as old as Daniel. Find their ages now.
:
Let D = Daniels age now; J = Justin's age now
:
Write an equation for each statement:
:
"In 4 years Daniel's age will be the same as Justin's age now."
D + 4 = J
:
"In 2 years,Justin will be twice as old as Daniel."
J + 2 = 2(D + 2)
J + 2 = 2D + 4
J = 2D + 4 - 2
J = 2D + 2
:
Find their ages now.
:
From the 1st statement; substitute (D+4) for J in the above equation, find D:
D + 4 = 2D + 2
4 - 2 = 2D - D
D = 2 yrs old now
Then
J = 2 + 4 = 6 yrs old now
:
:
Check our solution in the statement:
"In 2 years,Justin will be twice as old as Daniel."
6 + 2 = 2(2+2)
:
Did this make sense to you?